OpenStudy (anonymous):

x= -10 divided by -5 0ver 2

OpenStudy (anonymous):

can some one figure this out or help me

OpenStudy (raffle_snaffle):



OpenStudy (raffle_snaffle):

what this is from a pervious problem

OpenStudy (raffle_snaffle):

This is from the last problem we solved.

OpenStudy (ranga):

\[x = \frac{ -10 }{ \frac{ -5 }{ 2 } } = -10 \times \frac{ 2 }{ -5 } = 2 \times 2 = 4\]

OpenStudy (anonymous):

thank you so much

OpenStudy (ranga):

You are welcome.